# The Quantum Complexity of String Breaking in the Schwinger Model

### v1

@@ IV Results Obtained with Classical Simulations
where $\hat{Y}_{i_{k}}$ is the Pauli matrix acting on qubit $i_{k}$. One interesting $n$-tangle to examine is $\tau^{(2)}_{(i,i+1)}$ throughout the lattice, which provides a measure of nearest-neighbor entanglement.
Figure 4: Local and multipartite entanglement in string breaking. Top: the local vacuum-subtracted MI as a function of distance from the center $r$ and external charge separation $d$ for $N=880,a=0.25,g=0.09,m=0.04601$. The inset shows cross sections of the MI for a selection of $d$’s. Bottom: the vacuum-subtracted 2-tangle for the same parameters. The inset on the bottom panel shows the vacuum-subtracted 4-tangle. The black dashed lines indicate the peak of the charge distribution shown in Fig. 2. The red dashed lines correspond to the $d$ where the respective quantity changes sign for the smallest $r$.
The lower panel of Fig. 4 shows the vacuum-subtracted $\tau^{(2)}_{(i,i+1)}$ across the lattice as a function of the separation between the static charges. The nearest-neighbor entanglement is modified along the length of the string, which rapidly vanishes, changes sign in the vicinity of string breaking, then becomes confined into the mesons as they become isolated. This behavior is also evident in the vacuum-subtracted 4-tangle, $\tau^{(4)}_{(i,i+1,i+2,i+3)}-\tau^{(2)}_{(i,i+1)}\tau^{(2)}_{(i+2,i+3)}$, (inset of lower panel in Fig. 4) and in the vacuum-subtracted MI $\Delta I$, (upper panel of Fig. 4), both between adjacent spatial sites. The shape looks similar to $\tau^{(2)}_{(i,i+1)}$, indicating that much of, or all, the information is contained in the 2-tangle. [^10] We find the quantity $\tau^{(4)}_{(i,i+1,i+2,i+3)}-c\,\tau^{(2)}_{(i,i+1)}\tau^{(2)}_{(i+2,i+3)}\rightarrow 0$ for $c\approx 1.78$. This suggests that (local) entanglement structure is limited to physical sites. There is a separation for which $\tau^{(2)}_{(i,i+1)}$ vanishes along much of the string (indicated by the red dashed line), coinciding with the point of string breaking. [^11] We observe that the nonlocal 4-tangle i.e., $\tau^{(4)}_{(N/2,N/2+1,i,i+1)}-\tau^{(2)}_{(N/2,N/2+1)}\tau^{(2)}_{(i,i+1)}$} vanishes everywhere for intervals away from the center (i.e., except for nearest neighbors). Notably, the zero that develops in $\tau^{(2)}$ and the MI tracks the maximum of the charge distribution (shown with the black dashed line) after the string breaking. Deviations of $\Delta I$ from 0 show the localization of correlations as a function of $d$, reflecting the underlying structure of string breaking and meson formation. The $\Delta I$ in the middle region after string breaking arises from the nuclear force between the mesons, which vanishes when the mesons are sufficiently separated. It is interesting to note, that both the maximum and minimum of $\Delta I$ are within the width of the meson defined by the charge profile peak (Fig. 2).
@@
Figure 5: Internal structure of the outgoing meson revealed by the RoM. Top: the vacuum-subtracted RoM of adjacent sites, $\Delta R$ in units of $a$, shown for for $n\geq 440$ as a function of $d$, for $N=880,a=1/4$ as in Table 1. The inset shows the behavior of $\Delta R$ inside the meson region for $d=75$. Bottom: the charge density $q_{n}$ for $n\leq 440$ for the same system.
Figure 5 shows the vacuum-subtracted RoM, $\Delta R$ defined via Eq. (8) as
@@
of four adjacent lattice sites, i,,i+3i,...,i+3, as a function of position and the separation of the external charges. Here $R_{\Omega}$ denotes the value at $d=0$. With increasing $d$, the deviation of $\Delta R$ from zero increases along the string, becoming maximal at the point where the string breaks, then decreasing rapidly as the mesons separate. The (local) RoM reveals structure in the string state before it breaks, and in the outgoing mesons.
@@
The inset of Fig. 5 shows the distribution of $\Delta R$ within the meson at $d=75$. This suggests that moments of RoM distributions in hadronic structure calculations could be complementary probes of structure, in the same way that moments of the charge or axial distributions are computed (and measured experimentally). This shares similarities with the behavior of $\Delta I$ within the meson (inset of top panel in Fig. 4). The lattice-spacing dependence of $\Delta R$ is discussed in App. E. Compared to $\tau^{(2)}$, $\Delta R$ reflects the transition between strings and mesons less sharply.
Figure 6: Dependence of string breaking on the fermion mass. Left: $m/g=0$, center: $m/g=1/2$ and right: $m/g=1$. Top: the local vacuum-subtracted RoM as a function of $d$ over half of the lattice $n\geq N/2$ with $a=1$. Bottom: the associated charge densities for $n\leq N/2$ for the same parameters.
Figure 6 shows $\Delta R$ for different values of the fermion mass. The massless case has $\Delta R>0$ inside the isolated mesons, which is qualitatively different from the massive cases for which $\Delta R<0$. String breaking is found to happen more abruptly for larger $d$, which is naively expected due to the increasing energy associated with screening the static charges. Furthermore, no string breaking is observed in the massless case, a result of dynamical fermions being easily “pulled” from the vacuum to screen the external charges. A more complete analysis can be found in App. E.
@@
A novel class of quantum correlations in the context of string breaking are those between spatially-separated subregions [103]. The mutual information $I(A:B)$, defined in Eq. (6), provides a measure of both classical and quantum correlations between regions $A$ and $B$.
Figure 7: Mutual information in string breaking. Left: The vacuum-subtracted MI, $I(A:B)$, between the spatial site (two staggered sites) adjacent to the midpoint of the lattice and the spatial site of the vertical axis as a function of the separation between the static charges, obtained with simulation parameters described in the text using $N=880$ and $a=1/4$. The dashed line at $d=46.5$ indicates the peak of entanglement entropy, antiflatness and nonlocal magic, as shown in Fig. 3. Right: The corresponding vacuum-subtracted MI between the spatial sites of the static charges.
@@
The left panel of Fig. 7 shows $I(A:B)$ between spatial sites along the string with respect to the center of the system. The $\Delta I$ is continuous along the string, peaking at the point of string breaking, and rapidly returning to vacuum values thereafter. Perhaps it is not surprising that the string provides long-distance classical or quantum correlations, or both, along the length of the string. It is also unsurprising that these correlations vanish as the system transitions to isolated mesons. However, the transition region extends (at least) over the range of the nuclear force between the mesons. [^12] It would be interesting to examine the behavior of $I(A:B)$ and $\Delta R^{\text{NL}}_{AB}$ for a system with unnatural scattering parameters, to determine whether it is the string-breaking scale, the nuclear force or the scattering parameters that determine the change in the measures as a function of separation. The vacuum-subtracted MI over the interval between the two external charges, where region $A$ is fixed at the position of the first external charge, is shown in the right panel of Fig. 7. Analogous to the left panel, it is continuous along the string and its extent peaks near the string breaking. At larger $d$, it is nonzero over the extent of the meson at the reference static charge.
@@
Figure 8: Nonlocal quantum correlations in string breaking as measured by the NL RoM. Left: the RoM of disjoint regions, $\Delta R_{AB}$, where region $A$ is at the center of the lattice and region $B$ is distance $r$ away. The vacuum value ($d=0$) and the values as $r\to\infty$ are subtracted. The system parameters $N=220,a=1,m_{\text{lat}}=0.045,g=0.09$ are used as in Table 1. Right: the NL RoM, $\Delta R^{\text{NL}}_{AB}$ for the same system parameters.
@@
The left panel of Fig. 8 shows the RoM $\Delta R_{AB}$ between disjoint subsystems on the lattice as a function of the distance between the regions $r$. Region $A$ is at the center of the lattice, $A=(N/2+1,N/2+2)$, and region $B$ is distance $r$ away, $B=A+(r,r+1)$. We define this quantity with a different vacuum subtraction to that in Eq. (19),
@@
where $R_{\infty}$ denotes the quantity evaluated as $r\rightarrow\infty$ and $R_{\Omega}$ denotes the value at $d=0$ (as before). The reason for this is that the wavefunction is modified at the center of the lattice before string breaking, providing a modification in $R$ for all values of $i$. The isolated meson is visible in $\Delta R_{AB}$, as is the string breaking. Unlike the MI, the RoM shows a different structure along the string in the vicinity of the string breaking. In particular, it exhibits a localized rapid change of sign at the point of string breaking (as identified by the vanishing of the vacuum-subtracted 2-tangle and other measures). This is consistent with the center being the location of the most “restructuring” of the wavefunction. This rapid increase in charge density (also seen in Fig. 2) coincides with the “hot spot” in the left panel of Fig. 8. Interestingly, the NL RoM (defined in Eq. (10)) shows similar, but more localized behavior to the MI (cf. Fig. 7). The outgoing meson “track” vanishes in the NL components of the RoM (right panel of Fig. 8), suggesting that NL quantum information is only encoded in the string state. The presence of NL RoM establishes a link between spatially extended physical objects (strings) and nonlocal quantum complexity independent of basis choice. These purely quantum correlations, absent from traditional models of hadronization, could imprint themselves into final states in high-energy collisions.
V Summary and Discussion
Understanding confinement and developing robust predictive capabilities for the resulting dynamical phenomena, such as fragmentation and hadronization, remains a long-standing challenge. On the path to scientific discovery in high-energy and nuclear physics, techniques from quantum information science are now being used to address this challenge. We have performed classical simulations of string formation and breaking in the (Abelian) Schwinger model, focusing on quantum complexity beyond bipartite entanglement entropy. Measures of both local and nonlocal quantum complexity are found to provide distinct and complementary probes of the mechanism of string formation and breaking.
@@ V Summary and Discussion
We find that quantities, such as the nonstabilizerness and multipartite entanglement of contiguous subsystems, exhibit rapid variations in the vicinity of string breaking, in contrast to classical quantities like energy density. Importantly, nonlocal quantum complexity (as measured by both the MI and the NL RoM) is found along the string, with classical correlations appearing to be somewhat longer range than the quantum correlations. Both vanish abruptly after string breaking. This shows that both classical and quantum correlations develop during string breaking, but suggests that the longer distance correlations are predominately classical. Unfortunately, our computations of nonlocal measures of complexity are limited to small systems due to the resources required for minimizations in the subregions, along with considerations of the stabilizer polytope. We have not computed measures of higher-body nonlocal magic, for the reasons stated above, and cannot rule them out.
@@
The quantum complexity exhibits sharp changes in the transition from flux tubes to isolated hadronic bound states, akin to phase transitions [21]. We expect that the behavior of quantum complexity in a dynamical simulation of string breaking will help answer questions difficult to investigate with statics alone. For instance, the dynamics of charge extraction from the vacuum can shine light on the mechanisms governing confinement. Such a study would require the simulation of wavepacket evolution [70, 104, 105] or the use of a time-dependent Hamiltonian [106, 107]. Applying our techniques to dynamical settings, to non-Abelian theories and to higher-dimensional systems will be illuminating. We also anticipate that the quantum complexity in string breaking will be imprinted in the final states of collisions of hadrons and nuclei. Analyses are already underway at the LHC to quantify entanglement and magic in top-anti-top production events [108, 109]. If detected in experiment, the hierarchies of these correlations and their (non)locality may affect modeling of fragmentation and hadronization.
Acknowledgments
@@
We would like to thank Roland Farrell, Henry Froland, Tobias Haug, Dima Kharzeev, Eliana Marroquin and Caroline Robin for helpful discussions. This work was supported, in part, by U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science. S.G. was supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Grants No. DE-FG02-97ER-41014 and in part by a Feodor Lynen Research fellowship of the Alexander von Humboldt foundation. This work was also supported, in part, by the Department of Physics and the College of Arts and Sciences at the University of Washington. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award NP-ERCAP0032083. This research was done using services provided by the OSG Consortium [110, 111, 112, 113], which is supported by the National Science Foundation awards #2030508 and #1836650. We have made use of the ITensor library for tensor network computations [114, 115, 116].
@@ Appendix A Analysis of lattice spacing artifacts
Table 2: The values of $d$ for which the measures (left column) reach their maximum value, for parameters in Table 1 and both orientations of the external field $E_{\text{ext}}=$- and $E_{\text{ext}}=+$, and the average of the two orientations.
@@ Appendix B OBC vs PBC
With the parameters chosen as described in Sec. II.2, the effect of boundaries and finite volume on the physics of the simulation is exponentially suppressed. As a result, the choice of boundary conditions should not impact the results of the computations. [^13] OBCs are highly preferred due to the inefficiency of standard MPS in representing periodic states [117]. With OBCs, dynamical gauge field degrees of freedom can be eliminated from the theory, resulting in long range fermionic interactions through a linear potential. In PBCs, there is a single gauge field degree of freedom, the “zero mode”, whose dynamics is not constrained by Gauss’s law. We follow the approach of Refs. [105, 118, 119], where the gauge field zero mode is truncated to a single state and zero mode dynamics is not considered. This approximation is valid for low-energy dynamics, and corrections are suppressed by $L$. The Hamiltonian is modified from Eq. (1):
$\hat{H}_{\text{PBC}}\,(d)\$$=\ \frac{m_{\text{lat}}}{2}\sum_{n=1}^{N}(-1)^{n}\hat{Z}_{n}-\frac{g^{2}}{2}\sum_{n=1}^{N}\left\{\sum_{s=1}^{N_{\text{phys}}}\left(s-\frac{s^{2}}{N}\right)\left(1-\frac{\delta_{s,N_{\text{phys}}}}{2}\right)\hat{Q}_{n}\hat{Q}_{n+s}+2\sum_{k=1}^{N-1}\frac{N-k}{N}E_{\text{ext},n}(d)\hat{Q}_{n+k}\right\}$
$\quad+\frac{1}{2}\sum_{n=1}^{N-1}\left(\hat{X}_{n}\hat{X}_{n+1}+\hat{Y}_{n}\hat{Y}_{n+1}\right)+\frac{1}{2}(-1)^{N_{\text{phys}}+1}\left(\hat{X}_{N}\hat{X}_{1}+\hat{Y}_{N}\hat{Y}_{1}\right)\ .$ (21)
With the lattice parameters chosen as in Table 1, we observe a difference in the string breaking process between PBCs and OBCs with the same system size $N$. As shown in Fig. 10, the string breaks at a larger separation $d$ for PBCs. This is attributed to image charge effects in PBCs that are absent in OBCs, [^14] This is verified by considering a larger lattice with the same lattice spacing, where the image charges are further away and the string breaks at an $d$ closer to OBCs. indicating that these calculations are not far into regime prescribed in Sec II.2. The boundary effects, both from the lattice boundary in OBCs and image charges in PBCs, extend into the lattice volume by a length scale set by the lightest hadron mass. As expected, the charge density of the OBC and PBC vacua determined by DMRG agree in the bulk.
Figure 10: The charge density for OBC (left) and PBC (right) $N=220,a=1,m_{\text{lat}}=0.045,g=0.09$ as in Table 1.

### v2

@@ Abstract
where $\hat{Y}_{i_{k}}$ is the Pauli matrix acting on qubit $i_{k}$. Specifically, the 2-tangle $\tau^{(2)}_{(i,i+1)}$ provides a measure of nearest-neighbor entanglement.
Figure 3: Local and multipartite entanglement in string breaking. Top: the local vacuum-subtracted MI as a function of distance from the center $r$ and external charge separation $d$ for $N=880,a=1/4,g=0.09,m=0.04601$. The inset shows cross sections of the MI for a selection of $d$’s. Bottom: the vacuum-subtracted 2-tangle for the same parameters. The inset on the bottom panel shows the vacuum-subtracted 4-tangle. The black dashed lines indicate the peak of the charge distribution shown in Fig. 1. The red dashed lines show $d$ where the respective quantity changes sign for the smallest $r$.
The lower panel of Fig. 3 shows the vacuum-subtracted $\tau^{(2)}_{(i,i+1)}$ across the lattice as a function of $d$. The nearest-neighbor entanglement is modified along the length of the string, which rapidly vanishes, changes sign in the vicinity of string breaking, then becomes confined into the mesons as they become isolated. This behavior is also evident in the vacuum-subtracted 4-tangle, $\tau^{(4)}_{(i,i+1,i+2,i+3)}-\tau^{(2)}_{(i,i+1)}\tau^{(2)}_{(i+2,i+3)}$, (inset of lower panel in Fig. 3) and in the vacuum-subtracted MI $\Delta I$, (upper panel of Fig. 3), both between adjacent spatial sites. The shape looks similar to $\tau^{(2)}_{(i,i+1)}$, indicating that much of the information is contained in the 2-tangle. [^7] We find the quantity $\tau^{(4)}_{(i,i+1,i+2,i+3)}-c\,\tau^{(2)}_{(i,i+1)}\tau^{(2)}_{(i+2,i+3)}\rightarrow 0$} for $c\approx 1.78$. Similarly, the nonlocal 4-tangle ${\tau^{(4)}_{(N/2,N/2+1,i,i+1)}-\tau^{(2)}_{(N/2,N/2+1)}\tau^{(2)}_{(i,i+1)}}\to 0$ for $i$ away from the center. This suggests that (local) entanglement structure is limited to physical sites. There is a separation for which $\tau^{(2)}_{(i,i+1)}$ vanishes along much of the string (indicated by the red dashed line), coinciding with the point of string breaking. Notably, the zero that develops in $\tau^{(2)}$ and $\Delta I$ tracks the maximum of the charge distribution (shown with the black dashed line) after the string breaking. Deviations of $\Delta I$ from 0 show the localization of correlations as a function of $d$, reflecting the underlying structure of string breaking and meson formation. The $\Delta I$ in the central region after string breaking arises from the nuclear force between the mesons, which vanishes when the mesons are sufficiently separated. It is interesting to note, that both the maximum and minimum of $\Delta I$ are within the width of the meson defined by the charge profile peak (Fig. 1).
@@
Figure 4: Internal structure of the outgoing meson revealed by the RoM. Top: the vacuum-subtracted RoM of adjacent sites, $\Delta R$, in units of $a$, shown for $n\geq 440$ as a function of $d$, for $N=880,a=1/4,g=0.09,m=0.04601$. The inset shows $\Delta R$ inside the meson region for $d=75$. Bottom: the charge density $q_{n}$ for $n\leq 440$ for the same system. The black dashed lines show the external charge positions, and the red dashed lines mark the peaks of the bipartite complexity measures shown in Fig. 2.
Figure 4 shows the vacuum-subtracted RoM, $\Delta R$, defined via Eq. (13) as
@@
of four adjacent lattice sites, $i,\dots,i+3$, as a function of position and the separation of the external charges. Here $R_{\Omega}$ denotes the value at $d=0$. With increasing $d$, the deviation of $\Delta R$ from zero increases along the string, becoming maximal at the point where the string breaks, then decreasing rapidly as the mesons separate. The (local) RoM reveals structure in the string state before it breaks, and in the outgoing mesons.
@@
The inset of Fig. 4 shows $\Delta R$ within the meson at $d=75$. This suggests that moments of RoM distributions in hadronic structure calculations could be complementary probes of structure, in the same way that moments of the charge or axial distributions are computed (and measured experimentally). This shares similarities with the behavior of $\Delta I$ within the meson (inset of top panel in Fig. 3). The dependence of $\Delta R$ on lattice spacing and $m/g$ is discussed in App. E.
@@
A novel class of quantum correlations in the context of string breaking are those between spatially-separated subregions [105]. The mutual information $\Delta I(A:B)$ provides a measure of both classical and quantum correlations between regions $A$ and $B$.
Figure 5: Mutual information in string breaking. Left: The vacuum-subtracted MI, $\Delta I(A:B)$, where region $A$ is at the center of the lattice, $A=\left(\frac{N}{2}+1,\frac{N}{2}+2\right)$ and $B$ is a distance $r$ away, $B=A+(r,r+1)$. The parameters $N=880,a=1/4,g=0.09,m=0.04601$ are used. Right: $\Delta I(A:B)$ where $A$ is at the left external charge, $A=\left(\frac{N}{2}-\frac{d}{2a}-1,\frac{N}{2}-\frac{d}{2a}\right)$, and $B$ is a distance $r$ away, $B=A+(r,r+1)$. The dashed lines mark the peak of the bipartite complexity measures shown in Fig. 2.
@@
The left panel of Fig. 5 shows $\Delta I(A:B)$ between spatial sites along the string with respect to the center of the system. The MI is continuous along the string, peaking at the point of string breaking, and rapidly returning to vacuum values thereafter. Perhaps it is not surprising that the string supports long-distance correlations (classical or quantum) along the length of the string, which vanish as the system transitions to isolated mesons. However, the transition region extends (at least) over the range of the nuclear force between the mesons. [^8] The behavior of $\Delta I$ and $\Delta R^{\text{(NL)}}_{AB}$ for a system with unnatural scattering parameters may determine whether the string-breaking scale, the nuclear force or the scattering parameters drive changes in quantum complexity. The right panel of Fig. 5 shows $\Delta I$ over the interval between the static charges, where region $A$ is fixed at the position of the left charge. Its behavior along the string is analogous to the left panel, and it is nonzero over the extent of the meson for larger $d$.
@@
Figure 6: Nonlocal quantum correlations in string breaking as measured by the NL RoM. Left: the RoM of disjoint regions, $\Delta R_{AB}$, where region $A$ is at the center of the lattice and region $B$ is a distance $r$ away. The vacuum value ($d=0$) and the values as $r\to\infty$ are subtracted. The system parameters $N=220,a=1,m_{\text{lat}}=0.045,g=0.09$ are used. Right: the NL RoM, $\Delta R^{\text{(NL)}}_{AB}$ for the same system parameters. The dashed lines mark the peak of the bipartite complexity measures shown in Fig. 2.
@@
The left panel of Fig. 6 shows the RoM $\Delta R_{AB}$ between disjoint subsystems on the lattice as a function of the distance between the regions $r$. As in the left panel of Fig. 5, region $A$ is at the center of the lattice, A=(N/2+1,N/2+2){A=(N/2+1,N/2+2)}, and region $B$ is distance $r$ away, B=A+(r,r+1){B=A+(r,r+1)}. We define this quantity with a different vacuum subtraction to that in Eq. (18),
@@
where $R_{\infty}$ denotes the quantity evaluated as $r\rightarrow\infty$ and $R_{\Omega}$ denotes the value at $d=0$ (as before). The reason for this is that the wavefunction is modified at the center of the lattice before string breaking, providing a modification in $R$ for all values of $r$. Unlike $\Delta I$, $\Delta R_{AB}$ exhibits a localized rapid change of sign at the point of string breaking. This is consistent with the center being the location of the most “restructuring” of the wavefunction. This increase in charge density seen in Fig. 1 coincides with the “hot spot” in the left panel of Fig. 6. The behavior of $\Delta R_{AB}$ in the transition region immediately after string breaking may reveal new insights into the strong nuclear force. Interestingly, the NL RoM (defined in Eq. (15)) shows similar, but more localized behavior to the MI (cf. Fig. 5). The outgoing meson “track” vanishes in the NL components of the RoM (right panel of Fig. 6), suggesting that NL quantum information is only encoded in the string state. The presence of NL RoM establishes a link between spatially extended physical objects (strings) and nonlocal quantum complexity independent of basis choice. These purely quantum correlations, absent from traditional models of hadronization, could imprint themselves into final states in high-energy collisions.
Discussion— Continuing along the path of scientific discovery in high-energy and nuclear physics, techniques from quantum information science are now being used to understand confinement and develop predictive capabilities for the resulting dynamical phenomena.
@@
We have performed classical simulations of string formation and breaking in the (Abelian) Schwinger model, finding that quantum complexity varies rapidly during string breaking, and provides distinct probes complementary to classical quantities like energy density. Importantly, nonlocal quantum complexity (as measured by both the MI and the NL RoM) is found along the string. This shows that both classical and quantum correlations develop during string breaking, and suggests that the longer distance correlations are predominately classical. Our nonlocal complexity computations are limited to small systems due to the resources required for subregion minimizations and stabilizer polytope considerations. For these reasons, we have not computed higher-body nonlocal magic measures and cannot rule them out.
@@
The quantum complexity exhibits sharp changes in the transition from flux tubes to isolated hadronic bound states, akin to phase transitions [21]. Dynamical simulations of string breaking, via wavepacket evolution [70, 106, 107] or time-dependent Hamiltonians [108, 109], would clarify mechanisms of charge extraction and confinement that are difficult to investigate with statics alone. Applying our techniques to non-Abelian and higher-dimensional systems will be illuminating. We also anticipate that the quantum complexity in string breaking will be imprinted in final states of collisions of hadrons and nuclei. Analyses quantifying entanglement and magic in top-anti-top production events at the LHC are already underway [110, 111]. If detected in experiment, the hierarchies of these correlations and their (non)locality may affect modeling of fragmentation and hadronization.
@@
Acknowledgements—We would like to thank Roland Farrell, Henry Froland, Tobias Haug, Dima Kharzeev, Eliana Marroquin and Caroline Robin for helpful discussions. This work was supported, in part, by U.S. Department of Energy, Office of Science, Office of Nuclear Physics, InQubator for Quantum Simulation (IQuS) under Award Number DOE (NP) Award DE-SC0020970 via the program on Quantum Horizons: QIS Research and Innovation for Nuclear Science. S.G. was supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, Grants No. DE-FG02-97ER-41014 (UW Nuclear Theory) and in part by a Feodor Lynen Research fellowship of the Alexander von Humboldt foundation. This work was also supported, in part, by the Department of Physics and the College of Arts and Sciences at the University of Washington. This work was enabled, in part, by the use of advanced computational, storage and networking infrastructure provided by the Hyak supercomputer system at the University of Washington. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award NP-ERCAP0032083. This research was done using services provided by the OSG Consortium [112, 113, 114, 115], which is supported by the National Science Foundation awards #2030508 and #1836650. We have made use of the ITensor library for tensor network computations [116, 117, 118].
@@ Appendix A Analysis of lattice spacing artifacts
Table 1: The values of $d$ for which the measures (left column) reach their maximum value, for parameters in the main text and both orientations of the external field $E_{\text{ext}}=$- and $E_{\text{ext}}=+$, and the average of the two orientations.
@@ Appendix B OBC vs PBC
With the parameters chosen as described in the main text, the effect of boundaries and finite volume on the physics of the simulation is exponentially suppressed. As a result, the choice of boundary conditions should not impact the results of the computations. [^9] OBCs are highly preferred due to the inefficiency of standard MPS in representing periodic states [119]. With OBCs, dynamical gauge field degrees of freedom can be eliminated from the theory, resulting in long range fermionic interactions through a linear potential. In PBCs, there is a single gauge field degree of freedom, the “zero mode”, whose dynamics is not constrained by Gauss’s law. We follow the approach of Refs. [107, 120, 121], where the gauge field zero mode is truncated to a single state and zero mode dynamics is not considered. This approximation is valid for low-energy dynamics, and corrections are suppressed by $L$. The Hamiltonian is modified from Eq. (1):
$\hat{H}_{\text{PBC}}\,(d)\$$=\ \frac{m_{\text{lat}}}{2}\sum_{n=1}^{N}(-1)^{n}\hat{Z}_{n}-\frac{g^{2}a}{2}\sum_{n=1}^{N}\left\{\sum_{s=1}^{N_{\text{phys}}}\left(s-\frac{s^{2}}{N}\right)\left(1-\frac{\delta_{s,N_{\text{phys}}}}{2}\right)\hat{Q}_{n}\hat{Q}_{n+s}+\frac{2}{g}\sum_{k=1}^{N-1}\frac{N-k}{N}E_{\text{ext},n}(d)\hat{Q}_{n+k}\right\}$
$\quad+\frac{1}{4a}\sum_{n=1}^{N-1}\left(\hat{X}_{n}\hat{X}_{n+1}+\hat{Y}_{n}\hat{Y}_{n+1}\right)+\frac{1}{4a}(-1)^{N_{\text{phys}}+1}\left(\hat{X}_{N}\hat{X}_{1}+\hat{Y}_{N}\hat{Y}_{1}\right)\ .$ (20)
With the lattice parameters chosen as in the main text, we observe a difference in the string breaking process between PBCs and OBCs with the same system size $N$. As shown in Fig. 8, the string breaks at a larger separation $d$ for PBCs. This is attributed to image charge effects in PBCs that are absent in OBCs, [^10] This is verified by considering a larger lattice with the same lattice spacing, where the image charges are further away and the string breaks at an $d$ closer to OBCs. indicating that these calculations are not far into regime prescribed in the main texts. The boundary effects, both from the lattice boundary in OBCs and image charges in PBCs, extend into the lattice volume by a length scale set by the lightest hadron mass. As expected, the charge density of the OBC and PBC vacua determined by DMRG agree in the bulk.
Figure 8: The charge density for OBC (left) and PBC (right) $N=220,a=1,m_{\text{lat}}=0.045,g=0.09$.